OpenStudy (anonymous):

a farmer builds a pen against a stone wall. Fencing is used for the other 3 sides and is 13 ft long. The area is 15 ft sq. what are the dimensions?

OpenStudy (anonymous):

2 sides are 5 and one side is 3. 5+5+3=13 5*3=15 i solved by using a system of equations: 2x+y=13 and xy=15 x is one side and y is the other.

OpenStudy (anonymous):

so it is unnecessary to use the quadratic equation?

OpenStudy (anonymous):

no u could use it. This is wht i did: 2x+y=13 xy=15 2x+y=13 y=13-2x then substitute y into the othr equation. x(13-2x)=15 13x-2x^2=15 -2x^2+13x-15=0 you could use the quadratic formula.I chose to factor it.You will get the same answer. (-2x+3)(x-5) x=5 or 3/2 5y=15 y=3 or 3/2y=15 y=10 fina answr can be 10 , 3/2, 3/2 or 5, 5, 3

OpenStudy (anonymous):

makes since